# On the $\ell^{2}$-Betti numbers and algebraic fibring of the (outer) automorphism group of a right-angled Artin groups

### v2

@@ Abstract
We compute the first $\ell^{2}$-Betti number of the automorphism and outer automorphism groups of arbitrary right-angled Artin groups (RAAGs), providing a complete characterization of when it is non-zero. In addition, we determine all $\ell^{2}$-Betti numbers of the outer automorphism group in the case where the defining graph is disconnected and the associated RAAG is not isomorphic to a free group. We also analyse the algebraic fibring of the pure symmetric automorphism groups $\mathrm{PSA}(A_{\Gamma})$ and $\mathrm{PSO}(A_{\Gamma})$ and the virtual algebraic fibring of $\operatorname{Out}(A_{\Gamma})$ in the case when $A_{\Gamma}$ admits no non-inner partial conjugation. In the transvection-free case, we show that $\beta_{1}^{(2)}(\operatorname{Out}(A_{\Gamma}))=0$ if and only if $\operatorname{Out}(A_{\Gamma})$ virtually fibres.
@@ 1 Introduction
Regarding higher-dimensional $\ell^{2}$-Betti numbers, we obtain a complete description in the case of RAAGs defined by disconnected graphs, provided the RAAG is not isomorphic to a free group.
Proposition 1.3.
Let $\Gamma$ be a disconnected simplicial graph such that $A_{\Gamma}$ is not isomorphic to a free group. Then, all $\ell^{2}$-Betti numbers of $\operatorname{Out}(A_{\Gamma})$ vanish, except of the following case:
$\beta_{2}^{(2)}(\operatorname{Out}(\mathbb{Z}^{2}\ast\mathbb{Z}^{2}))=\frac{1}{2^{11}3^{2}}$
The higher $\ell^{2}$-Betti numbers of the outer automorphism groups of RAAGs remain largely mysterious when the defining graph $\Gamma$ is connected or when $A_{\Gamma}$ is isomorphic to a free group.
@@
When $A_{\Gamma}$ is transvection-free, i.e. if $u,v\in V(\Gamma)$ are such that $\operatorname{lk}(u)\subset\operatorname{st}(v)$ then $u=v$, the Torelli subgroup is of finite index in $\operatorname{Out}(A_{\Gamma})$. As a consequence of Theorems 1.2 and 1.4, we verify that Conjecture 1.5 holds for $\operatorname{Out}(A_{\Gamma})$ in this case.
@@
The paper is structured as follows. In section $2$, we present the necessary definitions and preliminary results that will be used throughout the text. Section $3$ is devoted to the study of virtual fibring, where we prove Theorems 1.4 and 1.7. In section $4$, we study the $\ell^{2}$-Betti numbers of $\operatorname{Aut}(A_{\Gamma})$ and $\operatorname{Out}(A_{\Gamma})$, proving Theorem 1.2 and Proposition 1.3. In the final section, we present several examples of RAAGs for which $\operatorname{Aut}(A_{\Gamma})$ and $\operatorname{Out}(A_{\Gamma})$ exhibit distinct fibring properties and $\ell^{2}$-Betti number behaviour.
@@ 2.1 $\ell^{2}$-Betti numbers
Follows from [25, Theorem 7.2 (1), (2), (7)]. ∎
@@
For $k=1$ the result follows from Lemma 2.1 For $k=2$ it is known that $F_{2}$ embeds into $\mathrm{GL}_{2}(\mathbb{Z})$ as a subgroup of index $24$ (c.f. [26]). Since $\beta_{1}^{(2)}(F_{2})=1$ and $\beta_{n}^{(2)}(F_{2})=0$ for all $n\neq 1$ the result follows from Proposition 2.2 For $k\geq 3$ the result follows from the fact that all latices in a given locally compact group are measure equivalent, that the vanishing of $\ell^{2}$-Betti numbers is invariant under measured equivalence and that $\beta_{n}^{(2)}(\mathrm{GL}_{k}(\mathbb{R}))=0$ for all $n\geq 0$ (c.f. [7] and [25] Theorem 7.34). ∎
@@ 2.2 Algebraic fibring
A group $G$ is said to fibre algebraically if there exists a non-trivial epimorphism $\varphi:G\to\mathbb{Z}$ with finitely generated kernel. By Proposition 2.3.3 it follows that if $G$ virtually fibres algebraically then $\beta_{1}^{(2)}(G)=0$.
@@
Let $G_{1}$ be a finite index subgroup of $G$ that fibres, with fibration $f:G_{1}\to\mathbb{Z}$. Consider the preimage $H_{1}=\varphi^{-1}(G_{1})$, which is a finite index subgroup of $H$. Let $\varphi_{1}:H_{1}\to G_{1}$ be the restriction of $\varphi$ to $H_{1}$. Then, $\ker(\varphi_{1})=\ker(\varphi)\cap H_{1}$, which is a finite index subgroup of the finitely generated group $\ker(\varphi)$. Hence, $\ker(\varphi_{1})$ is finitely generated. Now, since $G_{1}$ fibres and there is and there is an epimorphism $\varphi_{1}:H_{1}\to G_{1}$ with finitely generated kernel, it follows from Lemma 2.7 that $H_{1}$ is algebraically fibred. Therefore, $H$ virtually fibres. ∎
@@ 2.3 Right-angled Artin groups
Free groups and direct products of RAAGs are easy to understand in terms of the defining graph $\Gamma$: $A_{\Gamma}$ splits as a non trivial direct product $A_{\Gamma_{1}}\times A_{\Gamma_{2}}$ if and only if $\Gamma$ decomposes as a join $\Gamma_{1}\ast\Gamma_{2}$ and $A_{\Gamma}$ splits as a non trivial free product $A_{\Gamma_{1}}\ast A_{\Gamma_{2}}$ if and only if $\Gamma$ decomposes as a disjoint union $\Gamma_{1}\sqcup\Gamma_{2}$. Both facts follows from the isomorphism theorem for RAAGs, i.e. $A_{\Gamma}\cong A_{\Theta}$ if and only if $\Gamma\cong\Theta$. A proof of this result may be found in [24].
@@
If $A_{\Gamma}$ is virtually fibred then $\beta_{1}^{(2)}(A_{\Gamma})=0$ so, by Theorem 2.11, $\overline{\beta}_{0}(\widehat{\Gamma})=0$ so $\Gamma$ is connected. If $\Gamma$ is connected then $BB_{\Gamma}$ is finitely generated, so the map $\varphi:A_{\Gamma}\to\mathbb{Z}$ given by sending all standard generators to one is a fibration of $A_{\Gamma}$. Conversely, if $\Gamma$ is disconnected $\beta_{1}^{(2)}(A_{\Gamma})=\overline{\beta}_{0}(\widehat{\Gamma})\neq 0$ so $A_{\Gamma}$ is not virtually fibred. ∎
@@ 2.4 Automorphism group of a RAAG
  • Transvections: if $v,w\in V(\Gamma)$ the map $\tau^{v}_{w}$ given by $\tau^{v}_{w}(w)=wv$ and $\tau^{v}_{w}(u)=u$ for every $u\in V(\Gamma)\setminus\{u\}$ is a well defined element of $\operatorname{Aut}(A_{\Gamma})$ if and only if $w\leq v$.
@@
Let $\Gamma$ be a simplicial graph and $u,v\in V(\Gamma)$ be two vertices that are not connected in $\Gamma$. We say that the pair $(u,v)$ forms a separating intersection of links (SIL) if there is a connected component of $\Gamma\setminus(\operatorname{lk}(u)\cap\operatorname{lk}(v))$ that does not contain $u$ nor $v$.
@@
If the graph $\Gamma$ has at least $3$ connected components $\Gamma_{1},\Gamma_{2}$ and $\Gamma_{3}$ then for any two vertices $u\in V(\Gamma_{1})$ and $v\in V(\Gamma_{2})$, the pair $(u,v)$ forms a SIL. Indeed, since $\Gamma_{3}$ is a connected component of $\Gamma\setminus(\operatorname{lk}(u)\cap\operatorname{lk}(v))$, it witness the SIL condition. Hence, we may assume that $\Gamma=\Gamma_{1}\sqcup\Gamma_{2}$ with $\Gamma_{1}$ and $\Gamma_{2}$ connected graphs.
@@
To conclude this section, we present several results that will later serve as tools for constructing examples of RAAGs exhibiting distinct $\ell^{2}$-Betti numbers and diferent algebraic fibering properties.
@@ 2.5 Pure symmetric automorphism group
These explicit presentations can be found in [23]. In the case when the graph $\Gamma$ has no SILs the group $\mathrm{PSA}(A_{\Gamma})$ is isomorphic to a RAAG.
Theorem 2.23 (Charney-Ruane-Stambaugh-Vijayan, [9]).
If $\Gamma$ is a simplicial graph with no SILs then $\mathrm{PSA}(A_{\Gamma})$ is isomorphic to the RAAG $A_{\Theta}$, where $\Theta$ is a simplicial graph constructed as follows:
  • The vertex set $V(\Theta)$ is bijective to the set of partial conjugations in $\operatorname{Aut}(A_{\Gamma})$.
  • Two vertices of $\Theta$ are conected with an edge if and only if the corresponding partial conjugations commute in $\operatorname{Aut}(A_{\Gamma})$.
Corollary 2.24.
If $\Gamma$ is a simplicial graph such that $\Gamma\setminus\operatorname{st}(v)$ is connected for every $v\in V(\Gamma)$ then $\Gamma$ has no SILs and $\mathrm{PSA}(A_{\Gamma})$ is isomorphic to $A_{\Gamma}$.
@@
For any pair of vertices $u,v\in V(\Gamma)$ we have $\Gamma\setminus\operatorname{st}(v)\subset\Gamma\setminus(\operatorname{lk}(u)\cap\operatorname{lk}(v))$. Since $\Gamma\setminus\operatorname{st}(v)$ is connected by assumption, it follows that $\Gamma\setminus(\operatorname{lk}(u)\cap\operatorname{lk}(v))$ is also connected. Therefore, $A_{\Gamma}$ is a RAAG with no SILs. By Theorem 2.23, $\mathrm{PSA}(A_{\Gamma})$ is isomorphic to the RAAG $A_{\Theta}$, where the vertex set is in bijection with the set of inner automorphisms $\{\iota_{v}\}_{v\in V(\Gamma)}$ of $A_{\Gamma}$. Since two inner automorphism $\iota_{v}$ and $\iota_{w}$ commute if and only if the corresponding generators $v$ and $w$ commute in $A_{\Gamma}$, i.e. $\{v,w\}\in E(\Gamma)$, it follows that $\Theta\cong\Gamma$. ∎
@@
The graph $\Theta$ is given the following edges:
@@ 3.1 Pure symmetric automorphism group
Therefore, we may assume that $\Gamma\setminus\operatorname{st}(v)$ is connected for every $v\in V(\Gamma)$. Hence, by Corollary 2.24, $\Gamma$ has no SILs and $\mathrm{PSA}(A_{\Gamma})$ is isomorphic to $A_{\Gamma}$. By Lemma 2.17 either $A_{\Gamma}$ is connected or $A_{\Gamma}\cong\mathbb{Z}^{n}\ast\mathbb{Z}^{m}$ for some $m,n\geq 1$. In the latter case, it follows by Theorem 4.15.ii in [21] that $\beta_{1}^{(2)}(\mathbb{Z}^{n}\ast\mathbb{Z}^{m})=1$ and thus $A_{\Gamma}$ does not virtually fibre. We may thus assume that $\Gamma$ is connected. In this case $A_{\Gamma}$ fibres due to Corollary 2.14. ∎
@@
and $\chi(\pi^{u}_{C})=0$ for any other partial conjugation $\pi^{u}_{C}$. The map $\chi$ is well-defined map because all the relators in the standard presentation of $\mathrm{PSO}(A_{\Gamma})$ are commutators except for the relators of the form $\prod_{C\in I_{w}}\pi^{w}_{C}=1$ where $w\in V(\Gamma)$ and $I_{w}$ is the set of connected components of $\Gamma\setminus\operatorname{st}(v)$. These relations are clearly preserved under $\chi$. The support of $\chi$ equals to $\{\pi_{L}^{v},\pi_{K}^{v},\pi^{u}_{C}\}$ which consists of $3$ partial conjugations based at the same vertex $v$, but corresponding to three different connected components of $\Gamma\setminus\operatorname{st}(v)$. Therefore, the support of $\chi$ is not the subset of a $\delta$-p-set and, by Theorem 3.5.2, $\chi]\in\Sigma^{1}(\mathrm{PSO}(A_{\Gamma}))$. Consequently, $\mathrm{PSO}(A_{\Gamma})$ fibres. ∎

### v3

@@ Abstract
We compute the first $\ell^{2}$-Betti number of the automorphism and outer automorphism groups of arbitrary right-angled Artin groups (RAAGs), providing a complete characterization of when it is non-zero. We also analyse the algebraic fibring of the pure symmetric automorphism groups $\mathrm{PSA}(A_{\Gamma})$ and $\mathrm{PSO}(A_{\Gamma})$ and the virtual algebraic fibring of $\operatorname{Out}(A_{\Gamma})$ in the case when $A_{\Gamma}$ admits no non-inner partial conjugation. In the transvection-free case, we show that $\beta_{1}^{(2)}(\operatorname{Out}(A_{\Gamma}))=0$ if and only if $\operatorname{Out}(A_{\Gamma})$ virtually fibres.
@@ 1 Introduction
The higher $\ell^{2}$-Betti numbers of the (outer) automorphism groups of RAAGs remain largely mysterious in general.
@@
When $A_{\Gamma}$ is transvection-free, i.e. if $u,v\in V(\Gamma)$ are such that $\operatorname{lk}(u)\subset\operatorname{st}(v)$ then $u=v$, the Torelli subgroup is of finite index in $\operatorname{Out}(A_{\Gamma})$. As a consequence of Theorems 1.2 and 1.3, we verify that Conjecture 1.4 holds for $\operatorname{Out}(A_{\Gamma})$ in this case.
@@
The paper is structured as follows. In section $2$, we present the necessary definitions and preliminary results that will be used throughout the text. Section $3$ is devoted to the study of virtual fibring, where we prove Theorems 1.3 and 1.6. In section $4$, we study the $\ell^{2}$-Betti numbers of $\operatorname{Aut}(A_{\Gamma})$ and $\operatorname{Out}(A_{\Gamma})$, proving Theorem 1.2. In the final section, we present several examples of RAAGs for which $\operatorname{Aut}(A_{\Gamma})$ and $\operatorname{Out}(A_{\Gamma})$ exhibit distinct fibring properties and $\ell^{2}$-Betti number behaviour.
@@ 2.1 $\ell^{2}$-Betti numbers
Follows from [24, Theorem 7.2 (1), (2), (7)]. ∎
@@
An important well-known fact that we will use is that the first $\ell^{2}$-Betti number of a finitely generated group is always finite. Since we could not find a proof in the literature stated, we include a short argument. The proof relies only on standard properties of $\ell^{2}$-homology and von Neumann dimension (cf. [24]).
Lemma 2.5.
Let $G$ be an infinite finitely generated group. Then $\beta_{1}^{(2)}(G)\leq d(G)-1,$ where $d(G)$ denotes the minimal number of generators of $G$.
Proof.
If $d=d(G)$, we may construct a $K(G,1)$-complex by taking a bouquet of $d$ circles and attaching higher-dimensional cells. Passing to the universal cover, we obtain a cellular chain complex of Hilbert $\mathcal{N}(G)$-modules of the form
$\cdots\xrightarrow{\partial_{2}}\ell^{2}(G)^{d}\xrightarrow{\partial_{1}}\ell^{2}(G)\to 0.$
The first $\ell^{2}$-Betti number is given by $\beta_{1}^{(2)}(G)=\dim_{\mathcal{N}(G)}(\ker\partial_{1})-\dim_{\mathcal{N}(G)}(\overline{\mathrm{Im}(\partial_{2})})$. Using the additivity of the von Neumann dimension in weakly exact sequences (cf. [24, Theorem 1.12(2)]), we have
$\dim_{\mathcal{N}(G)}(\ker\partial_{1})=\dim_{\mathcal{N}(G)}(\ell^{2}(G)^{d})-\dim_{\mathcal{N}(G)}(\overline{\mathrm{Im}(\partial_{1})}).$
Substituting, we obtain
$\beta_{1}^{(2)}(G)=\dim_{\mathcal{N}(G)}(\ell^{2}(G)^{d})-\dim_{\mathcal{N}(G)}(\overline{\mathrm{Im}(\partial_{1})})-\dim_{\mathcal{N}(G)}(\overline{\mathrm{Im}(\partial_{2})}).$
Since $G$ is infinite, the zeroth $\ell^{2}$-Betti number vanishes by Lemma 2.1, so the complex is exact at $\ell^{2}(G)$. Thus
$\dim_{\mathcal{N}(G)}(\overline{\mathrm{Im}(\partial_{1})})=\dim_{\mathcal{N}(G)}(\ell^{2}(G))=1.$
Finally, $\beta_{1}^{(2)}(G)=d-1-\dim_{\mathcal{N}(G)}(\overline{\mathrm{Im}(\partial_{2})})\leq d-1$, as required. ∎
@@
For $k=1$ the result follows from Lemma 2.1 For $k=2$ it is known that $F_{2}$ embeds into $\mathrm{GL}_{2}(\mathbb{Z})$ as a subgroup of index $24$ (c.f. [25]). Since $\beta_{1}^{(2)}(F_{2})=1$ and $\beta_{n}^{(2)}(F_{2})=0$ for all $n\neq 1$ the result follows from Proposition 2.2. For $k\geq 3$ the result follows from the fact that all latices in a given locally compact group are measure equivalent, that the vanishing of $\ell^{2}$-Betti numbers is invariant under measured equivalence (c.f. [17]) and that $\beta_{n}^{(2)}(\mathrm{GL}_{k}(\mathbb{R}))=0$ for all $n\geq 0$ (c.f. [7] and [24] Theorem 7.34). ∎
@@ 2.2 Algebraic fibring
A group $G$ is said to fibre algebraically if there exists a non-trivial epimorphism $\varphi:G\to\mathbb{Z}$ with finitely generated kernel. By Proposition 2.3 it follows that if $G$ virtually fibres algebraically then $\beta_{1}^{(2)}(G)=0$.
@@
Let $G_{1}$ be a finite index subgroup of $G$ that fibres, with fibration $f:G_{1}\to\mathbb{Z}$. Consider the preimage $H_{1}=\varphi^{-1}(G_{1})$, which is a finite index subgroup of $H$. Let $\varphi_{1}:H_{1}\to G_{1}$ be the restriction of $\varphi$ to $H_{1}$. Then, $\ker(\varphi_{1})=\ker(\varphi)\cap H_{1}$, which is a finite index subgroup of the finitely generated group $\ker(\varphi)$. Hence, $\ker(\varphi_{1})$ is finitely generated. Now, since $G_{1}$ fibres and there is an epimorphism $\varphi_{1}:H_{1}\to G_{1}$ with finitely generated kernel, it follows from Lemma 2.8 that $H_{1}$ is algebraically fibred. Therefore, $H$ virtually fibres. ∎
@@ 2.3 Right-angled Artin groups
Free products and direct products of RAAGs are easy to understand in terms of the defining graph $\Gamma$: $A_{\Gamma}$ splits as a non trivial direct product $A_{\Gamma_{1}}\times A_{\Gamma_{2}}$ if and only if $\Gamma$ decomposes as a join $\Gamma_{1}\ast\Gamma_{2}$ and $A_{\Gamma}$ splits as a non trivial free product $A_{\Gamma_{1}}\ast A_{\Gamma_{2}}$ if and only if $\Gamma$ decomposes as a disjoint union $\Gamma_{1}\sqcup\Gamma_{2}$. Both facts follows from the isomorphism theorem for RAAGs, i.e. $A_{\Gamma}\cong A_{\Theta}$ if and only if $\Gamma\cong\Theta$. A proof of this result may be found in [23].
@@
If $A_{\Gamma}$ is virtually fibred then $\beta_{1}^{(2)}(A_{\Gamma})=0$ so, by Theorem 2.12, $\overline{\beta}_{0}(\widehat{\Gamma})=0$ and $\Gamma$ is connected. If $\Gamma$ is connected then $BB_{\Gamma}$ is finitely generated, so the map $\varphi:A_{\Gamma}\to\mathbb{Z}$ given by sending all standard generators to one is a fibration of $A_{\Gamma}$. Conversely, if $\Gamma$ is disconnected $\beta_{1}^{(2)}(A_{\Gamma})=\overline{\beta}_{0}(\widehat{\Gamma})\neq 0$ so $A_{\Gamma}$ is not virtually fibred. ∎
@@ 2.4 Automorphism group of a RAAG
  • Transvections: if $v,w\in V(\Gamma)$ the map $\tau^{v}_{w}$ given by $\tau^{v}_{w}(w)=wv$ and $\tau^{v}_{w}(u)=u$ for every $u\in V(\Gamma)\setminus\{w\}$ is a well defined element of $\operatorname{Aut}(A_{\Gamma})$ if and only if $w\leq v$.
@@
Let $\Gamma$ be a simplicial graph and $u,v\in V(\Gamma)$ be two vertices that are not adjacent in $\Gamma$. We say that the pair $(u,v)$ forms a separating intersection of links (SIL) if there is a connected component of $\Gamma\setminus(\operatorname{lk}(u)\cap\operatorname{lk}(v))$ that does not contain $u$ nor $v$.
@@
If the graph $\Gamma$ has at least $3$ connected components $\Gamma_{1},\Gamma_{2}$ and $\Gamma_{3}$ then for any two vertices $u\in V(\Gamma_{1})$ and $v\in V(\Gamma_{2})$, the pair $(u,v)$ forms a SIL. Indeed, since $\Gamma_{3}$ is a connected component of $\Gamma\setminus(\operatorname{lk}(u)\cap\operatorname{lk}(v))$, it witnesses the SIL condition. Hence, we may assume that $\Gamma=\Gamma_{1}\sqcup\Gamma_{2}$ with $\Gamma_{1}$ and $\Gamma_{2}$ connected graphs.
@@
To conclude this section, we present several results that will later serve as tools for constructing examples of RAAGs exhibiting distinct $\ell^{2}$-Betti numbers and different algebraic fibring properties.
@@ 2.5 Pure symmetric automorphism group
These explicit presentations can be found in [22]. Before moving to the $\operatorname{Out}(A_{\Gamma})$ case, let us record a simple observation about $\mathrm{PSA}(A_{\Gamma})$.
Lemma 2.24.
Assume that $\Gamma$ is a simplicial graph such that $\Gamma\setminus\operatorname{st}(v)$ is connected for every $v\in V(\Gamma)$. Then $\Gamma$ has no SILs, and $\mathrm{PSA}(A_{\Gamma})$ is isomorphic to $A_{\Gamma\setminus\Delta}$, where $\Delta$ is the induced subgraph of $\Gamma$ spanned by those vertices $v\in V(\Gamma)$ with $\operatorname{st}(v)=\Gamma$.
@@
For any pair of vertices $u,v\in V(\Gamma)$ we have $\Gamma\setminus\operatorname{st}(v)\subseteq\Gamma\setminus(\operatorname{lk}(u)\cap\operatorname{lk}(v))$. Since $\Gamma\setminus\operatorname{st}(v)$ is connected by assumption, it follows that $\Gamma\setminus(\operatorname{lk}(u)\cap\operatorname{lk}(v))$ is also connected. Hence $\Gamma$ has no SILs.
On the other hand, every partial conjugation is an inner automorphism. Therefore, $\mathrm{PSA}(A_{\Gamma})=\mathrm{Inn}(A_{\Gamma})=A_{\Gamma}/Z(A_{\Gamma})$, and the result follows from Proposition 2.11. ∎
@@
The graph $\Theta$ is given by the following edges:
@@ 3.1 Pure symmetric automorphism group
Therefore, we may assume that $\Gamma\setminus\operatorname{st}(v)$ is connected for every $v\in V(\Gamma)$. By Corollary 2.24, $\Gamma$ has no SILs and $\mathrm{PSA}(A_{\Gamma})$ is isomorphic to $A_{\Gamma\setminus\Delta}$, where $\Delta$ is the subgraph of $\Gamma$ generated by the $v\in V(\Gamma)$ such that $\mathrm{st}(v)=\Gamma$. By Lemma 2.18 either $A_{\Gamma}$ is connected or $A_{\Gamma}\cong\mathbb{Z}^{n}\ast\mathbb{Z}^{m}$ for some $m,n\geq 1$. In the latter case, it follows by Theorem 4.15.ii in [20] that $\beta_{1}^{(2)}(\mathbb{Z}^{n}\ast\mathbb{Z}^{m})=1$ and thus $A_{\Gamma}$ does not virtually fibre. If $\Gamma$ is connected $A_{\Gamma}$ fibres due to Corollary 2.15. ∎
@@
and $\chi(\pi^{u}_{C})=0$ for any other partial conjugation $\pi^{u}_{C}$. The map $\chi$ is well-defined map because all the relators in the standard presentation of $\mathrm{PSO}(A_{\Gamma})$ are commutators except for the relators of the form $\prod_{C\in I_{w}}\pi^{w}_{C}=1$ where $w\in V(\Gamma)$ and $I_{w}$ is the set of connected components of $\Gamma\setminus\operatorname{st}(w)$. These relations are clearly preserved under $\chi$. The support of $\chi$ equals to $\{\pi_{L}^{v},\pi_{K}^{v},\pi^{u}_{C}\}$ which consists of $3$ partial conjugations based at the same vertex $v$, but corresponding to three different connected components of $\Gamma\setminus\operatorname{st}(v)$. Therefore, the support of $\chi$ is not the subset of a $\delta$-p-set and, by Theorem 3.5.2, $\chi]\in\Sigma^{1}(\mathrm{PSO}(A_{\Gamma}))$. Consequently, $\mathrm{PSO}(A_{\Gamma})$ fibres. ∎